Shawna bought some strings of beads. The number of strings was 7 more than the number of beads on each string. There were 60 beads in all. How many beads were on each string?
1
Expert's answer
2020-01-20T10:02:08-0500

Solution. Let x is the number of beads on each string, y is number of the strings. According to the condition of the problem, we have a system


"\begin{cases}
 x-y=7 \\
 x\times y=60
\end{cases}"

From the first equation of the system


"x=y+7"

Substituting into the second equation of the system we obtain


"x\times y=60 \implies (y+7)\times y=60"

"y^2+7y-60=0"

The roots of the quadratic equation


"y_1=5"

"y_2=-12<0"

Therefore the number of beads on each string is equal to 5.

Answer. 5.